OpenStudy (anonymous):

Find the radius of a circle with the given circumference. 6 square root of 2 pie in.

OpenStudy (anonymous):

C = 2 * pi * r \[6\sqrt{2} * \pi = 2 * \pi * r\] solve for r

OpenStudy (anonymous):

\[r=\sqrt{72}/2\]

OpenStudy (anonymous):

i forgot the pi add the pi in

OpenStudy (anonymous):

the pi's cancel out

OpenStudy (anonymous):

\[r = 3\sqrt{2}\]